OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true.

OpenStudy (anonymous):

(3d)3(d)

OpenStudy (anonymous):

(3d2)2

OpenStudy (anonymous):

The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(3d)^3 =???

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

you would cube each piece

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

what does cube it mean? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well something like 2 cubed would be 2*2*2 = 8 3 cubed = 3^3 = 3*3*3 = 27 4 cubed = 4*4*4 = 64 5 cubed = 5*5*5 = 125 etc etc

OpenStudy (anonymous):

so how many times do u want me to cube the 3? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you cube it once by multiplying three copies of 3

jimthompson5910 (jim_thompson5910):

3 cubed = 3^3 = 3*3*3 = 27

jimthompson5910 (jim_thompson5910):

and the other piece is d, so d^3 is just d^3

jimthompson5910 (jim_thompson5910):

(3d)^3 = 27d^3

OpenStudy (anonymous):

so now we just have to choose 1 of these

OpenStudy (anonymous):

The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

jimthompson5910 (jim_thompson5910):

(3d)^3 = 27d^3 (3d)^3*d = 27d^3*d (3d)^3*d = 27d^4

jimthompson5910 (jim_thompson5910):

the exponent in expression #1 is 4 do the same to find the exponent in expression #2

OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true.

OpenStudy (anonymous):

(3x2) 3x2

OpenStudy (anonymous):

(3x3)2(x2)

OpenStudy (anonymous):

The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

wait are we on a new one? or the same one?

OpenStudy (anonymous):

new

jimthompson5910 (jim_thompson5910):

oh so you figured out the other one then?

OpenStudy (anonymous):

i thought the answer was c.... is it?

jimthompson5910 (jim_thompson5910):

yes good, just checking

OpenStudy (anonymous):

okay now we can do the second 1

jimthompson5910 (jim_thompson5910):

as for the second one, please draw it out

OpenStudy (anonymous):

what do u mean draw it out?

jimthompson5910 (jim_thompson5910):

use the draw feature to show me the problem it's a bit vague

OpenStudy (anonymous):